codeine is a drug used in painkillers and is a weak base with kb 1.6X10-6. calculate the ph of a 2.2X10-3M solution of codeine. calculate % dissociation of codeine.

Question 1: Codeine (CiaH21 NO:) is a drug used in painkillers and is a weak base with a Kb value 1.6 x 106 1 solution of codeine (3 points) b) Calculate the % dissociation of codeine (2 points).
Answer #1

The Pkb of codeine = -Log(Kb) = -Log(1.6*10-6) = 5.8

a) pOH = 1/2 {pKb - Log[codeine]}

= 1/2 {5.8 - Log(2.2*10-3)}

= 4.23

Now, pH + pOH = 14

i.e. pH + 4.23 = 14

i.e. pH = 14 - 4.23 = 9.77

b) The percent dissociation of codeine = (Kb/[codeine])1/2 *100

= {(1.6*10-6)/(2.2*10-3)}*100

= 0.073%